OpenStudy (anonymous):

A prism has a right triangle as its base with a hypotenuse that measures 10 meters and legs that measure 6 meters and 8 meters. What is the surface area of this prism if the height of the prism is 8 meters?

OpenStudy (anonymous):

SA = 2 ( 6 * 8) + ( 6 + 8 + 10 ) * 8 = 96 + 192 = 288 m^2

OpenStudy (callisto):

surface area = 2(1/2)(6x8) + 10x8 +6x8 + 8x8 = 240 m^2